# Learning positional encodings in transformers depends on initialization

### v3

@@ 3.3 Learnable PE models discover ground truth attention maps and positions
Figure 3: A,B) We compared the attention maps of each learnable PE model to those derived from the “ground truth” 2d-fixed model. C) A strong rank correlation between generalization performance and the agreement of attention maps to the 2d-fixed model. D) We directly compared the learned PE embeddings to the ground truth 2d-fixed PE. Since the embedding dimensions of the learned PE models are random and not indexed in the same way to the 2d-fixed embedding, we matched their embedding dimensions using an orthogonal Procrustes transform. After matching the dimensions of the PE embeddings, we computed the L2 norm to calculate the distance between the learned PE and the ground truth 2d-fixed PE. E) The distance (L2 norm) between the learned PEs and 2d-fixed PE, for every $\sigma$. F) We found a strong relationship between the PE agreement with the ground truth PEs and generalization performance ($\rho=-0.977$). (See also Appendix Table A3.)
@@
Next, we sought to understand how different PE initializations influenced the learned representations (e.g., attention maps and learned PEs) within the transformer. Since the underlying structure of the LST paradigm is a 2D grid, we used the 2d-fixed model as the ground truth model. This allowed us to assess how learned attention maps would deviate from attention maps derived from optimal position information. First, we extracted the attention weights for each model, and computed the cosine similarity of attention weights between the ground truth and learnable PE models (Fig. 3A; Table A3). We then correlated the alignment of learned attention maps to the ground truth with generalization performance. We found that the degree of agreement of learned attention maps with the ground truth (2d-fixed) predicted improved generalization ($\rho=0.96$, $p<0.0001$; Fig. 3B). Importantly, learned PEs with small $\sigma$ tended to learn attention representations that were more aligned with the ground truth. (We also computed the Jensen-Shannon Divergence as a complementary distance measure, finding similar results; Appendix Fig. A9.) In the Appendix, we include comparative analyses of the attention maps of models with common PEs (1d-fixed, relative, etc.) and compare those to the ground truth model (Fig. A10B-D).
@@
Next, to directly interpret what embeddings the learned PEs converged to during training, we measured teh alignment of learned PEs with the ground truth 2d-fixed PEs (Fig. 3D). This involved estimating the distance (i.e., L2 norm) between the 2d-fixed PE and the learned PE embedding after an orthogonal Procrustes transform was applied. An orthogonal Procrustes transform was applied to rotate and match embedding dimensions according to maximal similarity since embedding dimensions were arbitrary in the learnable PE models. We found that small-norm initialized PEs could better approximate the PEs from the 2d-fixed PE scheme (Fig. 3E). Critically, more similar ground truth PE approximation (measured by L2 norm) near perfectly predicted downstream generalization ($\rho=-0.98$, $p<0.0001$; Fig. 3F). These results indicate that 1) low-norm initializations can discover ground truth PEs and their subsequent attention maps, and 2) these discovered PEs predicted downstream generalization.
4 Experiment 2: Masked prediction of a nonlinear stochastic network simulation
@@
In our next experiment, we sought to demonstrate the impact of transformer PEs in learning data representations from a nonlinear, stochastic network simulation. Unlike natural language text, many real-world networks, such as genetic networks, air traffic flows, and brain propagation networks, are inherently spatially embedded in physical or conceptual space. Thus, understanding the positional arrangements of embedded data – e.g., the spatial relation between two genes in DNA, or the functional relation between two brain regions – is of paramount importance when learning representations of that data. We first explore the importance of learning PEs in a toy network simulation embedded into clusters of networks, before subsequently investigating its importance in real-world brain data in the subsequent section.
4.1 Recovering network modules via PE learning in a stochastic network simulation
@@
We implemented a nonlinear multivariate autoregressive (NMAR) model to simulate a system of 15 nodes organized into 3 network clusters (5 nodes per network) (Fig 4A). Each node evolves over time based on a combination of: 1) Autoregressive effects from its own past values ($p=3$ lags); 2) Strong intra-cluster interactions with other nodes in the same cluster; 3) Weak inter-cluster interactions with nodes in different clusters; 4) Private noise to introduce variability.
The timeseries for node $x_{i}$ at time $t$ was computed as
$x_{i}(t)=\sum_{k=1}^{p}w_{i,k}\cdot x_{i}(t-k)+\sum_{j\in C_{i},j\neq i}\lambda_{ij}\cdot f(x_{j}(t-1))+\sum_{j\notin C_{i}}\eta_{ij}\cdot f(x_{j}(t-1))+\epsilon_{i}(t),$
where $w_{i,k}\sim\mathcal{U}(0.2,0.5)$, $C$ refers to the module, $\lambda_{ij}\sim\mathcal{U}(0.02,0.2)$, $\eta_{ij}\sim\mathcal{U}(0.005,0.01)$, $\epsilon_{i}(t)\sim\mathcal{N}(0,0.2)$, and $f(x)=\text{sin}(x)$.
@@
Figure 4: Recovering network structure in learned PEs using a simple nonlinear, stochastic network simulation. A) Network structure with 15 nodes divided into 3 modules, where intra-module nodes interacted more strongly with each other than inter-module nodes. B) Example time series from 3 randomly selected nodes. C) We used a self-supervised training objective, where the model was trained to predict masked data from contemporaneous timepoints. D) The network modularity of PEs, which was measured in relation to the “ground truth” networks, was highest in learnable PE models initialized with small norm (learn-0.1) (see Appendix A.3 for calculation). Boxplots are sorted by mean modularity in descending order. E) The equation governing the timeseries generation for node $x_{i}(t)$ (see text for description of parameters). F) The cosine similarity of the PEs for each pair of tokens, across all models. We can observe that learnable PE models with small-norm initializations (particularly learn-0.1) can recover the ground truth network structure, while others cannot.
@@
5 Experiment 3: Masked prediction of spontaneous brain imaging data
@@
In neuroscience, a central goal is to be able to predict distributed brain activity using the activity of other brain regions (Bassett and Sporns, 2017). To achieve this, we sought to build a generalizable transformer model that would predict the brain activity of target regions using the brain activity of other regions. In this context, we conceptualized ‘tokens’ as distinct brain regions across the cortical mantle with the goal of predicting masked brain activity. This task can be formalized as masked pretraining, where the input tokens to the model are contemporaneous brain activity across different brain regions with masked (or missing) activity values (Fig 5B). This requires the model to predict missing brain activity from its surrounding context, i.e., the brain activity values of other brain regions.
@@
Naively, the PE of brain regions might manifest as their physical location in 3D space. However, decades of neuroscience research has revealed a modular brain organization, whereby different brain regions belong to distinct functional networks (or communities) (Power et al., 2011; Yeo et al., 2011; Ji et al., 2019; Schaefer et al., 2018). Thus, the goal of using this real-world dataset was to assess the degree to which models could recover this “modular” network organization in terms of PE. We used publicly available human functional magnetic resonance imaging (fMRI) data from the Human Connectome Project (HCP) dataset (www.humanconnectomeproject.org). We used the resting-state fMRI data from a subset ($n=100$; $n_{train}=70$; $n_{test}=30$) of the HCP 1200 participant pool (Van Essen et al., 2013). fMRI data were partitioned into 360 distinct cortical regions (tokens), which mapped onto to 12 distinct functional networks (Ji et al., 2019; Glasser et al., 2016). Additional details of the fMRI data and preprocessing pipelines can be found in Appendix A.2.
5.1 Learning interpretable PEs in human brain data for generalized brain activity prediction
@@
We trained transformer models with a mix of fixed and learnable PEs (fixed: 1d-fixed, relative, rope, random; learnable PEs initialized with $\sigma\in\{0.1,0.2,1.0,2.0\}$), minimizing the MSE of masked brain activity. Figure 5 shows results of the training with 50% masking (Fig. 5C), and testing with 90% masking (Fig. 5D). (We show results with 15%, 75%, and 90% mask training in Appendix Figs. A12,A13,A14) Successful generalization of the trained model involved predicting the masked brain activity of a separate subject’s data (i.e., test subjects). As expected, we found that small-norm initialized PEs (learn-0.1 and learn-0.2) achieved the best performance in both their training and validation sets after a fixed number of training steps (50k). Interestingly, relative PEs fared the worst, followed by absolute 1d-fixed and random PEs.
Figure 5: A) The brain is organized into functional networks that are spatially distributed throughout the brain in 3D space. B) We sought to understand the influence of PE on a model’s ability to predict contemporaneous brain activity. This can be formalized as a masked prediction task, where a transformer is trained to predict contemporaneous, spatially masked brain activity. To test for generalization, we evaluated the MSE on a validation dataset, which involved predicting the masked activity of data collected from different participants with increased masking (90%). (Models were trained with a 50% masking. See Figures A12,A13,A14 for results with 15%, 75%, and 90% masked training objectives.) C) We found that transformers endowed with different PE schemes converged to different MSE loss values. D) Notably, learnable PEs initialized with small norms (learn-0.1 and learn-0.2) achieved to the lowest MSE for both training and validation datasets. (X-axis is sorted from highest to lowest MSE.)
@@
Having established the superior predictive performance of richly learned PEs (small $\sigma$), we next evaluated whether these models learned meaningful yet nontrivial position information from masked pretraining of brain activity. While the PE of brain regions could be encoded as their physical location in 3D space, decades of neuroscience research has revealed a modular brain network organization, whereby different brain regions that are spatially distributed throughout the cortex belong to distinct functional networks (or communities) (Fig. 5A) (Power et al., 2011; Yeo et al., 2011; Ji et al., 2019; Schaefer et al., 2018). In other words, two brain regions that are distant in 3D space may actually be “functionally” close (e.g., yellow regions in Fig. 5A). (Prior work has indicated this functional closeness is determined by anatomical connectivity; Vázquez-Rodríguez et al. (2019).) We therefore sought to address whether richly learned PEs could recover this modular functional organization by measuring the distance of learned PE parameters.
When flattening brain regions across the cortex into a 1D tensor, each brain region’s (i.e., token’s) assignments are distributed across that tensor (Fig. 6A). (In this context, a PE scheme that places adjacent tokens closer to each other, such as in the original 1d-fixed PE with sines and cosines from Vaswani et al. (2017), would be clearly ill-suited.) To evaluate whether learned PEs learned a modular organization that reflected the known functional network organization of the brain, we measured the distance between every pair of tokens’ PE. This involved computing an orthogonal Procrustes transform to rotate and match the embedding dimensions of each token’s PE prior to computing their distance (L2 norm) (Figure 6B). The reason this is necessary is because since learned PEs are first randomly initialized, the embedding dimension of each token’s PE are not necessarily aligned (e.g., position 1’s embedding dimension $i$ does not necessarily correspond to embedding dimension $i$ of position 2). After aligning the embedding dimensions across tokens, we computed the distance between every pair of tokens. (We then scaled distance in this 2D matrix to range from 0 and 1, and computed the complement (i.e., $1-d_{scaled}$), such that closer PEs would have higher values. We computed both the network modularity and network clustering (i.e., segregation) with respect to the known network partitions (see A.3 for mathematical definitions; Rubinov and Sporns (2010)). In brief, modularity is a statistic that quantifies the degree to which the distance matrix can be cleanly subdivided into the brain’s network partitions. Network clustering is a statistic that quantifies the ratio between within-module distances and across-module distances, where modules are defined using a network partitioning from Ji et al. (2019). We found that the modularity of small-norm initialized PEs (learn-0.1 and learn-0.2) had the highest overall network modularity and segregation relative to other learnable PEs. This implies that the small-norm initialized PEs learned interpretable PEs with respect to the known functional networks of the brain. These findings support the hypothesis that learnable PEs (as opposed to off-the-shelf PEs) in the rich training regime can improve brain activity prediction, while successfully learning interpretable position information (Figure 6C,D).
@@
Figure 6: A) We evaluated whether learnable PEs could discover the brain’s modular functional network organization. When flattening the brain from a 3D to 1D tensor, the network labeling of each brain region is heterogenously distributed across a 1D tensor in a disorganized manner. B) We computed the distance between positional embeddings between every pair of token positions after aligning their embedding dimensions ($d$) through an orthogonal Procrustes transform. This allowed us to construct a token-by-token embedding distance matrix, which we then compared to brain network organization. C) The modularity of PEs with respect to the brain’s network organization. Models that had learnable PEs initialized from small-norm distributions learned a modular PE organization that was consistent with previously identified brain networks (Ji et al., 2019) D) The network clustering of PEs, which assessed whether PEs of tokens that belong to the same network are closer in space than PEs that do not belong to the same network.
@@ 6 Discussion
Related work. Recent studies have revealed that the choice of PE can strongly influence transformer generalization (Li et al., 2024; Kazemnejad et al., 2023; Golovneva et al., 2024; Ruoss et al., 2023; McLeish et al., 2024; Shen et al., 2024; Csordás et al., 2021; Ontanon et al., 2022; Zhou et al., 2024; Zhang et al., 2024). However, most of these investigations have been limited to evaluating the generalizability of various PE schemes on 1D string-based tasks (e.g., sequence learning tasks for arithmetic, context-free grammars, or compositional tasks). In contrast, many important problems require the encoding of sequences that are not in 1D (e.g., Li et al. (2021)), and where position information is non-trivial or not known, which we investigate here. Additionally, work in deep learning theory has provided insight into the impact of model initialization on representation learning, yet focus primarily on simple neural networks rather than transformers (Chizat et al., 2020; Woodworth et al., 2020; Jacot et al., 2020; Kunin et al., 2024; Lippl and Stachenfeld, 2024). In this study, we apply insights from deep learning theory to transformer models to effectively learn (and improve generalization) to nontrivial sequence tasks, such as tasks requiring reasoning in 2D, or tasks in which ground truth position information is organized in higher dimensions.
@@
Limitations and future directions. We have demonstrated that learnable PEs initialized from small-norm distributions can 1) approximate the ground truth PE, and 2) outperform many commonly-used PEs. However, there remain several limitations of the present study which future studies can explore. First, though we consider the use of the LST (with a 2D organization) a strength of this study due to the visual interpretability of the paradigm’s positional information, it is unclear how well this approach will generalize to tasks with an arbitrary number of elements or tasks in which there are dynamic changes in the number of elements (e.g., length generalization problems in higher dimensions). In addition, due to the task-dependent nature of utilizing (or learning) optimal PEs, for some tasks and training objectives, such as generic next-token prediction or arithmetic (which is order invariant under addition), standard PE choices may be most appropriate (e.g., 1d-fixed or rndpe). However, for tasks in which establishing an underlying ordering and relation of tokens is crucial — such as reasoning tasks in 2D or tasks with complex network structures, as is common in biology — our results show that using small-norm initialized learnable PEs can be highly beneficial. Second, the current learned PE is limited insofar that the embedding is linearly superimposed on a given token (i.e., $token+pe$). While this makes it potentially difficult to generalize to more complex tasks, a natural future research direction would be to learn nonlinear PE embeddings that allow for PEs to be flexibly generated as a function of the token embedding, thereby learning token abstractions/types (i.e., $pe(token)$). This nonlinear formulation of PE as a function of the token embedding would, in theory, have significantly greater expressive ability, and potentially endow transformers with the ability to recognize more complex formal languages (Merrill et al., 2024). Finally, while we have empirically demonstrated the applicability of prior learning theory on neural network initialization to the choice of PE initialization, in practice, it remains unclear how the precise choice of small $\sigma$ will impact generalization performance in practice (e.g., dependency on task and architectural choices, such as number of layers, embedding dimension, etc.). It will be important for future theoretical work to more carefully characterize how the choice of $\sigma$ may be influenced by transformer architecture and optimizer.
@@
Conclusion. There are many tasks and problems in which it is difficult to know the ground truth arrangement of input sequences, or tasks in which commonly-used PEs are ill-suited. Examples of such problems include reasoning on parse trees and directed graphs (where node distances and relations are not preserved when flattening into a 1D sequence for transformer sequence processing), or inference on real-world biological datasets in which the ground truth structure is important for prediction yet difficult to know (e.g., the 3D neuroscience data explored here, or co-expression of genes in a DNA sequence based on 3D chromatin conformation) (Szabo et al., 2019; Ji et al., 2021). In this study, we sought to understand how to learn position information directly from data using insights from deep learning theory. In particular, we found that an optimally-learned PE 1) outperformed commonly-used PEs, 2) learned attention maps and PE embeddings that were closely aligned to ground truth position information, and 3) enhanced generalization performance. Critically, learning an optimal and interpretable PE depended on its initialization in a reasoning task, a nonlinear netowrk simulation, and a real-world biological dataset. We anticipate these results will spur future investigations into the importance and utility of learnable PEs for structured learning and generalization.
7 Ethics statement
All human brain data used in this study is publicly available archived data from the Human Connectome Project (www.humanconnectomeproject.org). All participants gave signed, informed consent in accordance with the protocol approved by the local institutional review board.
8 Code availability
Code and task environments associated with reproducing models, results, and figures in this study will be made publicly available.
Appendix A Appendix / supplemental material

### v4

@@ 3.3 Learnable PE models learn interpretable attention maps and PEs
Figure 3: A,B) We compared the attention maps of each learned PE model to those derived from the “ground truth” 2d-fixed model. C) A strong rank correlation between generalization performance and the agreement of attention maps to the 2d-fixed model. D) We directly compared the learned PE embeddings to the ground truth 2d-fixed PE. Since the embedding dimensions of the learned PE models are random and not indexed in the same way to the 2d-fixed embedding, we matched their embedding dimensions using an orthogonal Procrustes transform. After matching the dimensions of the PE embeddings, we computed the L2 norm to calculate the distance between the learned PE and the ground truth 2d-fixed PE. E) The L2 distance between the learned PEs and 2d-fixed PE, for every $\sigma$. F) We found a strong relationship between the PE agreement with the ground truth PEs and generalization performance ($\rho=-0.977$). (See also Appendix Table A3.)
@@
Next, we sought to understand how different PE initializations influenced the learned representations (e.g., attention maps and learned PEs) within the transformer. Since the underlying structure of the LST paradigm is a 2D grid, we used the 2d-fixed model as the ground truth model. This allowed us to assess how learned attention maps would deviate from attention maps derived from optimal position information. First, we extracted the attention weights for each model, and computed the cosine similarity of attention weights between the ground truth and learnable PE models (Fig. 3A; Table A3). We then correlated the alignment of learned attention maps to the ground truth with generalization performance. We found that the degree of agreement of learned attention maps with the ground truth (2d-fixed) predicted improved generalization ($\rho=0.96$, $p<0.0001$; Fig. 3B). Importantly, learned PEs with small $\sigma$ tended to learn attention representations that were more aligned with the ground truth. (We also computed the Jensen-Shannon Divergence as a complementary distance measure, finding similar results; Appendix Fig. A8.) In the Appendix, we include comparative analyses of the attention maps of models with common PEs (1d-fixed, 1d-relative, etc.) and compare those to the ground truth model (Fig. A9B-D).
@@
Next, to directly interpret what embeddings the learned PEs converged to during training, we measured the alignment of learned PEs with the ground truth 2d-fixed PEs (Fig. 3D). This involved estimating the distance (i.e., L2 norm) between the 2d-fixed PE and the learned PE embedding after an orthogonal Procrustes transform was applied. An orthogonal Procrustes transform was applied to rotate and match embedding dimensions according to maximal similarity, since embedding dimensions were arbitrary in the learnable PE models. We found that small-norm initialized PEs could better approximate the PEs from the 2d-fixed PE scheme (Fig. 3E). Critically, more similar ground truth PE approximation (measured by L2 norm) near perfectly predicted downstream generalization ($\rho=-0.98$, $p<0.0001$; Fig. 3F). These results indicate that 1) low-norm initializations can learn interpretable PEs and attention maps, and 2) these interpretable representations predict generalization.
4 Experiment 2: Nonlinear network simulation
@@
In our next experiment, we sought to demonstrate the impact of PE initialization in transformers in learning data representations from a nonlinear, stochastic network simulation. Unlike natural language text, many real-world networks, such as genetic networks, air traffic flows, and brain propagation networks, are inherently spatially embedded in physical space. Thus, understanding the positional arrangements of embedded data – e.g., the spatial relation between two genes in DNA, or the functional relation between two brain regions – is of paramount importance when learning representations of that data. We first explore the possibility of learning PEs in a network simulation embedded into network clusters, before investigating its applicability to real-world brain data in the subsequent section.
@@
We implemented a nonlinear multivariate autoregressive (NMAR) model to simulate a system of 15 nodes organized into 3 network clusters (5 nodes per network) (Fig 4A). Each node evolves over time based on: 1) Autoregressive effects from its own past values ($p=3$ lags); 2) Strong intra-cluster interactions with other nodes in the same cluster; 3) Weak inter-cluster interactions with nodes in different clusters; 4) Private noise to introduce variability (Fig. 4C).
@@
Figure 4: Recovering network structure in learned PEs using a simple nonlinear, stochastic network simulation. A) Network structure with 15 nodes divided into 3 modules, where intra-module nodes interacted more strongly with each other than inter-module nodes. B) Example time series from 3 randomly selected nodes. C) The governing equation for node timeseries, where $w_{i,k}\sim\mathcal{U}(0.2,0.5)$, $C$ refers to the module, $\lambda_{ij}\sim\mathcal{U}(0.02,0.2)$, $\eta_{ij}\sim\mathcal{U}(0.005,0.01)$, $\epsilon_{i}(t)\sim\mathcal{N}(0,0.2)$, and $f(x)=\text{sin}(x)$. D) We used a self-supervised training objective, where the model was trained to predict masked data from contemporaneous timepoints. E) The network modularity of PEs, which was measured in relation to the “ground truth” networks, was highest in learnable PE models initialized with small norm (learn-0.1) (see Appendix A.3 for calculation). Boxplots are sorted by mean modularity in descending order. F) The cosine similarity of the PEs for each pair of tokens, across all models. We can observe that learnable PE models with small-norm initializations (particularly learn-0.1) can recover the ground truth network structure, while others cannot. (For additional data on convergence, see Fig. A16.)
@@
5 Experiment 3: Brain imaging data
@@
In neuroscience, an important goal is to be able to predict distributed brain activity using the activity of other brain regions (Bassett and Sporns, 2017). To achieve this, we sought to build a generalizable transformer model that would predict the brain activity of target regions using the brain activity of other regions. In this context, we conceptualized ‘tokens’ as distinct brain regions across the cortical mantle with the goal of predicting masked brain activity. This task can be formalized as masked pretraining, where the input tokens to the model are contemporaneous brain activity across different brain regions with masked (or missing) activity values (Fig A10B). This requires the model to predict missing brain activity from its surrounding context, i.e., the brain activity values of other brain regions.
@@
Naively, the PE of brain regions might manifest as their physical location in 3D space. However, decades of neuroscience research has revealed a modular brain organization, whereby different brain regions belong to distinct functional networks (or communities) (Power et al., 2011; Yeo et al., 2011; Ji et al., 2019; Schaefer et al., 2018). Thus, the goal of using this real-world dataset was to assess the degree to which models could recover this “modular” network organization in terms of PE. We used publicly available human functional magnetic resonance imaging (fMRI) data from the Human Connectome Project (HCP) dataset. We used the resting-state fMRI data from a subset ($n=100$; $n_{train}=70$; $n_{test}=30$) of the HCP 1200 participant pool (Van Essen et al., 2013). fMRI data were partitioned into 360 distinct cortical regions (tokens), which mapped onto to 12 distinct functional networks (Ji et al., 2019; Glasser et al., 2016). Additional details of the fMRI data and preprocessing pipelines can be found in Appendix A.2.
@@
We trained transformer models with a mix of fixed and learnable PEs (fixed: 1d-fixed, 1d-relative, 1d-rope, random; learnable PEs initialized with $\sigma\in\{0.1,0.2,1.0,2.0\}$), minimizing the MSE of masked brain activity. As expected, we found that small-norm initialized PEs (learn-0.1 and learn-0.2) achieved the best generalization performance (see Fig. A10, A11, A12, A13 for experiments with different masking levels). Interestingly, relative PEs fared the worst, followed by absolute 1d-fixed and random PEs.
@@
Having established the superior generalization performance of richly learned PEs (small $\sigma$), we next evaluated whether these models learned interpretable position information from masked pretraining of brain activity. While the PE of brain regions could be encoded as their physical location in 3D space, decades of neuroscience research has revealed a modular brain network organization, whereby different brain regions that are spatially distributed throughout the cortex belong to distinct functional networks (i.e., colored regions in Fig. 5A) (Power et al., 2011; Yeo et al., 2011). In other words, two brain regions that are distant in 3D space may actually be “functionally” close (e.g., yellow regions in Fig. 5A). (Prior work has indicated this functional closeness is determined by anatomical connectivity; Vázquez-Rodríguez et al. (2019).) We therefore sought to address whether richly learned PEs could recover this modular functional organization by measuring the distance of learned PE parameters.
@@
Figure 5: A) We evaluated whether learned PEs could discover the brain’s modular functional networks. When flattening the brain from a 3D to 1D tensor, the network labeling of each brain region is heterogenously distributed across a 1D tensor. B) We computed the distance between positional embeddings between every pair of token positions after aligning their embedding dimensions ($d$) through an orthogonal Procrustes transform. This allowed us to construct a token-by-token embedding distance matrix, which we then compared to known brain networks. C) The modularity and D) network clustering of PEs with respect to known brain networks (Ji et al., 2019). Models that had learnable PEs initialized from small-norm distributions learned a modular/clustered PE organization that was consistent with previously identified brain networks.
When flattening brain regions across the cortex into a 1D tensor, each brain region’s (i.e., token’s) assignments are distributed across that tensor (Fig. 5A). (In this context, a PE scheme that places adjacent tokens closer to each other, such as in the original 1d-fixed PE with sines and cosines from Vaswani et al. (2017), would be clearly ill-suited.) To evaluate whether learned PEs learned a modular organization that reflected the known functional network organization of the brain, we measured the distance between every pair of tokens’ PE. This involved computing an orthogonal Procrustes transform to rotate and match the embedding dimensions of each token’s PE prior to computing their distance (L2 norm) (Figure 5B). The reason this is necessary is because since learned PEs are first randomly initialized, the embedding dimension of each token’s PE are not necessarily aligned (e.g., position 1’s embedding dimension $i$ does not necessarily correspond to embedding dimension $i$ of position 2). After aligning the embedding dimensions across tokens, we computed the distance between every pair of tokens. (We then scaled distance in this 2D matrix to range from 0 and 1, and computed the complement (i.e., $1-d_{scaled}$), such that closer PEs would have higher values. We computed both the network modularity and network clustering with respect to the known network partitions (see A.3 for mathematical definitions). In brief, modularity is a statistic that quantifies the degree to which the distance matrix can be cleanly subdivided into the brain’s network partitions. Network clustering is a statistic that quantifies the ratio between within-network distances and between-network distances, where networks are defined using a network partitioning from Ji et al. (2019). We found that the modularity of small-norm initialized PEs (learn-0.1 and learn-0.2) had the highest overall network modularity and clustering relative to other learnable PEs. This implies that the small-norm initialized PEs learned interpretable PEs with respect to the known functional networks of the brain. These findings support the hypothesis that learnable PEs (as opposed to off-the-shelf PEs) in the rich training regime can improve brain activity prediction, while successfully learning interpretable position information (Figure 5C,D).
@@ 6 Discussion
Related work. Recent studies have revealed that the choice of PE can strongly influence transformer generalization (Li et al., 2024; Kazemnejad et al., 2023; Golovneva et al., 2024; Ruoss et al., 2023; McLeish et al., 2024; Shen et al., 2024; Csordás et al., 2021; Ontanon et al., 2022; Zhou et al., 2024; Zhang et al., 2024). However, most of these investigations have been limited to evaluating the generalizability of various PE schemes on 1D string-based tasks (e.g., arithmetic and/or compositional tasks). In contrast, many important problems require the encoding of sequences that are not in 1D (e.g., Li et al. (2021)), and where position information is non-trivial or not known, which we investigate here. Additionally, work in deep learning theory has provided insight into the impact of model initialization on representation learning, yet focus primarily on simple neural networks rather than transformers (Chizat et al., 2020; Woodworth et al., 2020; Jacot et al., 2020; Kunin et al., 2024; Lippl and Stachenfeld, 2024). In this study, we apply insights from deep learning theory to transformer models to learn interpretable representations for nontrivial sequence tasks.
@@
Limitations and future directions. We have demonstrated that learnable PEs initialized from small-norm distributions can 1) learn interpretable PEs, and 2) outperform many standard PEs. However, there remain several limitations of the present study which future studies can explore. First, though we consider the use of the LST (with a 2D organization) a strength of this study due to the visual interpretability of the paradigm’s positional information, it is unclear how well this approach will generalize to tasks with an arbitrary number of elements or tasks in which there are dynamic changes in the number of elements (e.g., length generalization problems in 2D or higher). In addition, due to the task-dependent nature of utilizing (or learning) optimal PEs, for some tasks and training objectives, such as generic next-token prediction or arithmetic (which is order invariant under addition), standard PE choices may be most appropriate (e.g., 1d-fixed or rndpe). However, for tasks in which establishing an underlying ordering and relation of tokens is crucial — such as reasoning tasks in 2D or tasks with complex network structures, as is common in biology — our results show that using small-norm initialized learnable PEs can be highly beneficial. Second, the current learned PE is limited insofar that the embedding is linearly superimposed on a given token (i.e., $token+pe$). While this makes it potentially difficult to generalize to more complex tasks, a natural future research direction would be to learn nonlinear PE embeddings that allow for PEs to be flexibly generated as a function of the token embedding, thereby learning token abstractions/types (i.e., $pe(token)$). This nonlinear formulation of PE as a function of the token embedding would, in theory, have significantly greater expressive ability. Finally, while we have empirically demonstrated that PE initialization impacts the learned representation and generalization in transformers, in practice, it remains unclear how the precise choice of $\sigma$ interacts with other hyperparameters (e.g., architectural choices and optimization protocols). For example, while theory alone suggests small norm initializations should produce rich representations, in practice, such initializations can be hard to train due to small gradients (Chizat et al., 2020). This can be empirically observed by the experiments we performed when using $\sigma$=0.01 and $\sigma$=0.05 (Fig. A4), where we find slower training convergence with smaller $\sigma$ using SGD. Interestingly, however, when Adam is used, $\sigma=0.01$ leads faster training convergence yet poor generalization, which is likely due to Adam’s adaptive learning rates. Since the learning rate in Adam is scaled by the magnitude of the gradient flow in Adam, this behavior likely produces large changes to the weights (during the initial small gradient updates) that pushes learning regime back into the kernel learning regime. It will be important for future theoretical work to better understand how the choice of $\sigma$, architecture, and optimizer interact to influence representation learning.
@@
Conclusion. There are many problems in which it is difficult to know the ground truth arrangement of input sequences. Examples of such problems include reasoning on parse trees and directed graphs (where node distances are not preserved when flattened to 1D), or inference on real-world biological datasets in which the ground truth structure is important for prediction yet difficult to know (e.g., the 3D neural data or co-expression of genes in a DNA sequence based on 3D chromatin conformation) (Szabo et al., 2019; Ji et al., 2021). In this study, we sought to understand how to learn position information directly from data using insights from deep learning theory. We found that an optimally-learned PE 1) outperformed standard PEs, 2) learned interpretable attention maps and PE embeddings, and 3) enhanced generalization performance. Critically, learning an interpretable PE depended on its initialization.
We hope these results spur further investigations into the importance and utility of learnable PEs for interpretable learning and generalization.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning through a perspective of interpretability. There are many potential societal consequences of such work, none which we feel must be specifically highlighted here.
Appendix A Appendix